1. Form the partial differential equation by eliminating
the arbitrary constants a and b from z=ax+by.
2. Form the partial differential equation by eliminating
the arbitrary constants a and b from z=ax
1
Expert's answer
2021-08-24T18:58:41-0400

"\text{1. Given \( z = ax + by\) then we will have the following: } \\
\displaystyle \frac{\partial z}{\partial x} = a \text{ and } \displaystyle \frac{\partial z}{\partial y} = b\\
\text{Inserting these into \(z = ax + by\) we have}\\
\displaystyle z = x\frac{\partial z}{\partial x} +y\frac{\partial z}{\partial y} \\
\text{which is equivalent to: } \\
x\frac{\partial z}{\partial x} +y\frac{\partial z}{\partial y} - z = 0 \\
\text{2. Given \( z = ax \) then we will have the following: } \\
\displaystyle \frac{\partial z}{\partial x} = a \\
\text{Inserting these into \(z = ax\) we have}\\
\displaystyle z = x\frac{\partial z}{\partial x}"